reduce the following quadratic form 3x²-2y²-z²+12yz-4xy+8zx to canonical form by an orthogonal transformation. Also find the rank, index signature and nature


1
Expert's answer
2021-03-24T12:34:58-0400

CORRECTED SOLUTION

The quadratic form 3x²-2y²-z²+12yz-4xy+8zx has the matrix

"A=\begin{pmatrix}
 3 & -2 & 4\\
 -2 & -2 & 6\\
4 &6 & -1
\end{pmatrix}"

Its characteristic polynomial is

"\det \begin{pmatrix}
 3-\lambda & -2 & 4\\
 -2 & -2-\lambda & 6\\
4 &6 & -1-\lambda
\end{pmatrix}=-\lambda^3 +63\lambda-162"

One can notice that "\lambda=3" is a root of this polynomial.

"\frac{\lambda^3-63\lambda+162}{\lambda-3}=\lambda^2+3\lambda-54=(\lambda+9)(\lambda-6)"

Therefore, "\lambda_1=3", "\lambda_2=6", "\lambda_3=-9" are all the eigenvalues of the given matrix.

Find the corresponding eigenvectors by solving the equations "AX=\lambda_iX" with i=1,2,3.

1) Case "\lambda_1=3". The equation "AX=3X" implies

"-2y+4z=0"

"-2x-5y+6z=0"

The fundamental solution of this system is "z=t, y=2t, x=-2t".

Therefore, the first eigenvector is "e_1=(-2,2,1)^T". "|e_1|=\sqrt{(-2)^2+2^2+1^2}=3"

2) Case "\lambda_2=6". The equation "AX=6X" implies

"-3x-2y+4z=0"

"-2x-8y+6z=0"

The fundamental solution of this system is "z=\begin{vmatrix}
 -3 & -2 \\
 -2 & -8
\end{vmatrix}t=20t, y=\begin{vmatrix}
 4 & -3 \\
 6 & -2
\end{vmatrix}t=10t, x=\begin{vmatrix}
 -2 & 4 \\
 -8 & 6
\end{vmatrix}t=-20t".

Therefore, the second eigenvector is "e_2=(-2,1,2)^T". "|e_2|=\sqrt{(-2)^2+1^2+2^2}=3"

3) Case "\lambda_3=-9". The equation "AX=-9X" implies

"12x-2y+4z=0"

"-2x+7y+6z=0"

The fundamental solution of this system is

"z=\begin{vmatrix}
 12 & -2 \\
 -2 & 7
\end{vmatrix}t=-80t, y=\begin{vmatrix}
 4 & 12 \\
 6 & -2
\end{vmatrix}t=-80t, x=\begin{vmatrix}
 -2 & 4 \\
 7 & 6
\end{vmatrix}t=-40t".

Therefore, the third eigenvector is "e_3=(1,2,2)^T". "|e_3|=\sqrt{1^2+2^2+2^2}=3"

In the orthonormal basis "e_1\/3, e_2\/3, e_3\/3" the quadratic form will be

"3y_1^2+6y_2^2-9y_3^2"

This is the normal form of the given quadratic form. From it one can see that the quadratic form is non-degenerate, indefinite, its rank=3, its index (the number of positive terms) =2 and its signature (the difference between the number of positive terms and the number of negative terms) is 1.